# Drainage and lifetime of thin liquid films: the role of salinity and convective evaporation

### v1

@@ 1 Introduction
Numerous studies have investigated the behavior of surface bubbles, with the goal of finding a relationship between the bubble size and the size distribution of the ejected drops (Lhuissier and Villermaux,, 2012; Anguelova and Huq,, 2017; Poulain et al.,, 2018; Berny et al.,, 2021; Shaw and Deike,, 2024; Jiang et al.,, 2022, 2024; Deike et al.,, 2022). Several pathways for drop formation have been identified depending on bubble size with associated distributions, however several questions remain for the complete modeling of the drop generation over the ocean: first, the role of salt on the drainage or lifetime of surface bubbles is still uncertain. It is not accounted for in the models but has a visible effect on bubble bursting as demonstrated by several lab experiments (Mårtensson et al.,, 2003; Park et al.,, 2014; May et al.,, 2016; Zinke et al.,, 2022; Dubitsky et al.,, 2023).
@@
An additional difficulty in the study of surface bubbles is the link between drainage and lifetime. For many of the drops ejected, the typical size is given by the thickness of the cap at the instant of burst (Lhuissier and Villermaux,, 2012; Poulain et al.,, 2018; Shaw and Deike,, 2024) such that both information (thickness over time and time of burst) are important. However, while drainage is deterministic and can be modeled as a function of a few parameters, the lifetime of surface bubbles is intrinsically stochastic and exhibits wide distributions. Poulain et al., (2018) discussed in their introduction the difficulty of obtaining these distributions even in laboratory settings as seemingly without changing experimental parameters, the lifetime can change drastically. Finding a link between drainage and lifetime fluctuations would therefore improve our understanding of those bubbles.
@@
Finally, the experiments described above have all been performed in a still (closed box) or uncontrolled atmosphere. Yet, the air surrounding the bubble plays a role in the drainage through the humidity (Poulain et al.,, 2018) and in the bursting process through the air density (Jiang et al.,, 2022). In realistic oceanic conditions, the bubbles move at the surface and are subjected to the wind: airflow over the cap of the bubbles may change the drainage laws or the lifetime distribution through several mechanisms. First, it may entrain the film (Burgess et al.,, 1999), changing the flow structure inside of the bubble cap and therefore altering drainage. Second, airflow can change the evaporative rate, by forcing the convection and replacing the natural convection that takes place around surface bubbles in quiescent environments (Dietrich et al.,, 2016; Boulogne and Dollet,, 2018).
@@
We set out to investigate the effect of water salinity, atmospheric humidity and airflow on thin films to identify how these parameters affect the drainage and burst. Here we start by considering flat soap films instead of the cap of surface bubbles as a model object to investigate the effect of humidity and salinity. Soap films present the advantage of a better controlled generation and geometry allowing for a very repeatable initial condition. They also present the benefit of a direct visualization of the motion in the film at all times through interferometry. In contrast to surface bubbles, soap films drain due gravity only (instead of curvature induced pressure gradient), and require large surfactant concentrations to be stable for a few seconds. As a consequence of this large surfactant concentration, soap films usually rupture near the top (Saulnier et al.,, 2014; Pasquet et al.,, 2024), in contrast to low contamination bubbles, which can burst anywhere in the cap (Champougny et al.,, 2016).
@@ 2 Experimental setup
The experiments are designed to generate soap films from a solution bath and to allow them to drain until they spontaneously burst. The films are held up by a frame made out of three wires (Nylon, diameter 100 µm) and their bottom border is connected to the bath (the detailed setup is shown in Fig. 1 (a)). The wires are attached to a 3D printed rigid frame and the films have a width of 29 mm and a height of 20 mm. The rigid frame is attached to a linear motor, allowing for a precisely controlled and repeatable generation. The films are pulled out of the solution with a constant velocity of 1 cm/s and brought to a stop gradually at the desired height. After generation, the films are held in place until they spontaneously burst: this event is automatically detected by the camera and the linear motor the lowers the frame back in the bath. After a short delay of 5 s, a new experiment can start.
@@
The solutions used to generate soap films were prepared with deionized water, adding NaCl in amounts ranging from 1 g/L to 147 g/L. A non-ionic surfactant (Triton X-100) was added with a concentration of three times the Critical Micelle Concentration (CMC) to stabilize the film. This surfactant was chosen to minimize the interactions between the dissolved ions and the surfactant. The static surface tension of the various solutions was measured using a Langmuir trough and was found be essentially independent of the salinity and equal to 32 mN/m (see App. A). At high concentrations, solutions of NaCl have a viscosity up to 2 times that of pure water (Kestin et al.,, 1981), to isolate the effect of viscosity, we therefore also tested solutions of 5 and 10% Glycerol by weight with the same surfactant. A summary of the experimental conditions is provided in Tab 1 and is separated in two parts: in the drainage part, the experiments are repeated five to ten times, and we explore many different conditions. Note that we also performed experiments with a different concentration of Triton X-100 (1 CMC) and with C12E9, another non-ionic surfactant. In the lifetime section, we focus on obtaining robust statistics in the zero salinity case, but changing the relative humidity.
@@
In order to explore the effect of mixing of the air in the chamber, we use Rayleigh-Bénard convection. We use a heating mat at the bottom of the chamber and Peltier plates at the top to achieve the desired temperature difference. The temperatures at the top and bottom surface are measured and controlled using analog sensors (Texas Instruments LM35DZ/LFT4), and the temperature at the level of the soap film is also measured to ensure that it remains between 21 and 22C. The heating mat and Peltier plates are duty-cycle controlled such that the wall temperature is constant. The Rayleigh number controlling the instability is $Ra=\rho_{a}\beta_{a}\Delta T\ell^{3}g/\mu_{a}\alpha_{a}=\mathcal{O}(10^{7})$, with $\ell=60$ cm the height of the chamber, $\beta$ the thermal expansion coefficient, $\rho_{a}$ the density of air, $\mu_{a}$ the viscosity of air, $\alpha_{a}$ the thermal diffusivity of air, and a temperature difference $\Delta T$. At this Rayleigh number, the flow in the chamber is turbulent (Drazin,, 2002) and the air surrounding the film is therefore homogeneous in space as any slow or local variations of the experimental conditions are suppressed. The instability sets in air in the chamber into motion, and the typical root-mean-square velocity is 5 mm/s (measured using Particle Image Velocimetry). We detail the effect of mixing the air in the chamber in Sec. 5; in the drainage section (Sec. 4) there is no mixing, but we show in Fig. 5 that convection does not alter the drainage significantly.
@@ 3 Qualitative observations
From moderate salinities and at any humidity, a new instability is occurring near the top of the film. There, thick patches are formed at the interface between the colored film and the black film (brown plumes at the top of Fig. 2 (b) and yellow area in panel (a)). These patches are thicker than the ones surrounding them and therefore fall down in the film because of buoyancy until they reach the level that matches their thickness. This instability, reminiscent of a Rayleigh-Taylor instability, occurs in many of the cases studied: with other surfactants and in cases where glycerol was used instead of salt. It is therefore not specific to salted films and most likely simply a feature of long-lasting films and was previously studied in different contexts (Seimiya and Seimiya,, 2021).
@@ 4 Drainage
Using the videos of the film draining and the interferometry technique described in Sec. 2 we are able to measure the thickness of any given point, over time. A typical drainage curve for all the cases we explore in shown in Fig. 3 (a): the thickness is a few microns and drains over a timescale ranging from a few tens of seconds to minutes. We can also recover vertical thickness profiles by interpolating between consecutive interference fringes, at a given instant. This technique works from a few seconds after the film has stopped being pulled out ($t\gtrsim 2$ s) and outside the regions where marginal regeneration patches are present ($z\lesssim 15$ mm). This allows us to investigate how the vertical profile of thickness evolves over time (illustrated in Fig. 3 (b)). The thickness increases almost linearly when moving towards the bottom of the film, with a slope $\partial h/\partial z$ that increases in time.
@@
Within the parameter range tested, film drainage is remarkably similar: without any rescaling the drainage curves almost collapse onto a single curve (Fig. 3 (c)). In particular the fact that the drainage curves for different humidities are the same (different markers) shows that in the regime of interest with films lasting at most a few hundred seconds, evaporation has a negligible role in the thinning. This property (which was previously noted in the case of continuously drawn films Champougny et al., (2018)) holds true regardless of salinity, despite the known effect of salt to alter evaporation rates (El-Dessouky et al.,, 2002). The effect of salinity (different colors), is to slightly shift the curves without altering the initial thickness of the film. The rate at which the film drains is therefore altered by salinity, but this change remains moderate in the conditions tested. Similarly, increasing the viscosity (orange and pink markers) slightly shifts the curves and slows down drainage. At very long times, the curves corresponding to Glycerol and NaCl solutions start to deviate from each other. We interpret this result as being due to the qualitatively different features visible in the films. The collapse of the drainage is robust to a change of surfactant and / or surfactant concentration in the cases we explored (not shown on Fig. 3).
At long times (after about 10 s), the drainage curve exhibit a power law behavior in time with an exponent close to -1. This exponent is coherent with previous experiments on soap films where evaporation does not play a major role and where the main draining mechanism is the exchange of patches of fluid with the Plateau borders at the sides of the film (Berg et al.,, 2005; Monier et al.,, 2024). These films are said to be mobile (Mysels,, 1964) in opposition to rigid films where the interface is contaminated to the point where it applies a no-slip boundary condition to the interstitial fluid. In that case, the equation describing the thinning of the film is (Pasquet et al.,, 2024):
$\frac{\partial h}{\partial t}=-\frac{\rho g}{4\mu}h^{2}\frac{\partial h}{\partial z},$ (1)
with $\mu$ and $\rho$ the viscosity and density of the liquid, respectively. A downwards Poiseuille flow develops inside of the film, leading to a thickness varying like $h\sim\sqrt{z/t}$, clearly not compatible with the present experiments. The surfactant gradient stabilizing the film can instead be taken into account by introducing a finite stress at the interface. This stress a priori depends on the surfactant used and on the surfactant concentration and should be computed by introducing the equation of state for the surface excess of surfactant (Champougny et al.,, 2015). Under the simplifying assumption of a constant slip length (or stress) at the interface, previous authors have numerically recovered a thinning law close to $t^{-1}$ (Berg et al.,, 2005).
@@
Figure 4: (a) Dimensionless thickness as a function of $z/H$, color indicates time.  (b) Thickness as a function of the rescaled parameter $(z/H)/(t/t_{0})$. The data shown in (a) and (b) corresponds to a film with a salinity of 72 g/L and a relative humidity of 70 %. The black line is a linear fit with slope 0.26.  (c) Thickness as a function of the rescaled parameter $(z/H)/(t/t_{0})$, for all salinities and relative humidities tested. Colors and markers are identical to Fig. 3: markers indicate relative humidity and colors indicate salinity or Glycerol concentration. Each marker corresponds to an average between 5 to 10 repetitions. (Inset) Measured proportionality constant as a function of the salinity (Eq. (2)). Errorbars represent the dispersion between repetitions.
@@
with $h_{\rm ref}=7$ µm, the thickness at generation and $t_{0}$ is the viscous drainage time. This typical timescale can be obtained by considering the Eq. (1):
$t_{0}=\frac{\mu H}{\rho gh_{\rm ref}^{2}}.$ (3)
@@
For the solutions considered here this typical viscous time is 40 to 50 s depending on the salinity. Using the drainage scaling in combination with the viscous timescale, we can see that the thickness data for all solutions reasonably collapses onto a straight line (Fig. 4 (c)). The slope $\mathcal{A}$ of this linear relationship is an order $\mathcal{O}(1)$ non-dimensional constant, which can be evaluated and is independent on the salinity ($\mathcal{A}\approx 0.3$, inset) or on the viscosity, when the viscous timescale varies by 20% over the same range. There is a small visible offset between different solutions that we attribute to small differences in the location of the origin $z=0$ between runs.
@@
The drainage scaling presented here is robust for all conditions tested of salinity and relative humidity but also to a change of surfactant concentration or surfactant (Triton X-100 at 1 CMC or C12E9 at 3 CMC). This law is only valid from a few seconds after the start of the drainage (Fig. 3 (c)) as it diverges for $t\ll t_{0}$. It also applies only in the bulk of the film, away from edge effects: marginal regeneration at the bottom or sides of the film and the boundary between the colored and black film at the top. As noted above, if the film was rigid instead of mobile (following the distinction from Mysels, (1964)), the thickness should follow $h\propto\sqrt{z/t}$ and this drainage scaling would break down.
For all the solutions tested in this study, the drainage follows the same scaling and only weakly depends on the parameters of the experiment (relative humidity and salinity). For salinities ranging from zero to four times that of seawater, the drainage rate is altered by at most 20%, only through viscous effects. We only notice deviations from the universal drainage scaling for very long times, comparing solutions rich in Glycerol against highly salty solutions or for the largest salinity. For solutions with moderate salinities (up to seawater concentration), this study of film drainage therefore allows us to conclude that (i) the only effect of NaCl on thin film drainage is through viscosity and (ii) relative humidity (and therefore evaporation) does not play a significant role in the drainage, regardless of salinity.
@@
This last remark should be contrasted with the very strong effect that relative humidity has on film lifetime. For a given solution, the lifetime can range from only a few seconds at 30% humidity to tens of minutes at with an almost saturated atmosphere at 90% humidity. The behavior seen here for soap films should be contrasted with the behavior of surface bubbles where evaporation can create Marangoni flows towards the cap, therefore slowing down drainage (this effect is however only significant if the atmosphere is heated) (Poulain et al.,, 2018; Chandran Suja et al.,, 2018). In the following we therefore focus on the effect of relative humidity on the lifetime of the films and obtain a scaling law explaining this dependence.

### v2

@@ 1 Introduction
Numerous studies have investigated the behavior of surface bubbles, with the goal of finding a relationship between the bubble size and the size distribution of the ejected drops (Lhuissier and Villermaux,, 2012; Anguelova and Huq,, 2017; Poulain et al.,, 2018; Berny et al.,, 2021; Shaw and Deike,, 2024; Jiang et al.,, 2022, 2024; Deike et al.,, 2022). Several pathways for drop formation have been identified depending on bubble size with associated distributions, however several questions remain for the complete modeling of the drop generation over the ocean: first, the role of salt on the drainage or lifetime of surface bubbles is still uncertain. It is not accounted for in the models but has a visible effect on bubble bursting as demonstrated by several lab experiments (Mårtensson et al.,, 2003; Park et al.,, 2014; May et al.,, 2016; Zinke et al.,, 2022; Dubitsky et al.,, 2023; Mazzatenta et al.,, 2024).
@@
An additional difficulty in the study of surface bubbles is the link between drainage and lifetime, i.e. identifying a single rupture criterion. For many of the drops ejected, the typical size is given by the thickness of the cap at the instant of burst (Lhuissier and Villermaux,, 2012; Poulain et al.,, 2018; Shaw and Deike,, 2024) such that both information (thickness over time and time of burst) are important. However, while drainage is deterministic and can be modeled as a function of a few parameters, the lifetime of surface bubbles is intrinsically stochastic and exhibits wide distributions. Poulain et al., (2018) discussed in their introduction the difficulty of obtaining these distributions even in laboratory settings as seemingly without changing experimental parameters, the lifetime can change drastically. Finding a link between drainage and lifetime fluctuations would therefore improve our understanding of those bubbles.
@@
Finally, the experiments described above have all been performed in a still (closed box) or uncontrolled atmosphere. Yet, the air surrounding the bubble plays a role in the drainage through the humidity (Poulain et al.,, 2018; Pasquet et al.,, 2022) and in the bursting process through the air density (Jiang et al.,, 2022). In realistic oceanic conditions, the bubbles move at the surface and are subjected to the wind: airflow over the cap of the bubbles may change the drainage laws or the lifetime distribution through several mechanisms. First, it may entrain the film (Burgess et al.,, 1999), changing the flow structure inside of the bubble cap and therefore altering drainage. Second, airflow can change the evaporative rate, by forcing the convection and replacing the natural convection that takes place around surface bubbles in quiescent environments (Dietrich et al.,, 2016; Boulogne and Dollet,, 2018).
@@
We set out to investigate the effect of water salinity, atmospheric humidity and airflow on thin films to identify how these parameters affect the drainage and burst. Here we start by considering flat soap films instead of the cap of surface bubbles as a model object to investigate the effect of humidity and salinity. Soap films present the advantage of a better controlled generation and geometry allowing for a very repeatable initial condition. They also present the benefit of a direct visualization of the motion in the film at all times through interferometry. The films however require large surfactant concentrations to be stable for a few seconds. As a consequence of this large surfactant concentration, soap films usually rupture near the top (Saulnier et al.,, 2014; Pasquet et al.,, 2024), in contrast to low contamination bubbles, which can burst anywhere in the cap (Champougny et al.,, 2016).
@@ 2 Experimental setup
The experiments are designed to generate soap films from a solution bath and to allow them to drain until they spontaneously burst. The films are held up by a frame made out of three wires (Nylon, diameter 100 µm) and their bottom border is connected to the bath (the detailed setup is shown in Fig. 1 (a)). The wires are attached to a 3D printed rigid frame and the films have a width of 29 mm and a height of 20 mm. The rigid frame is attached to a linear motor, allowing for a precisely controlled and repeatable generation. The films are pulled out of the solution with a constant velocity of 1 cm/s and brought to a stop gradually (using a constant deceleration of -2.5 cm/s2) at the desired height. After generation, the films are held in place until they spontaneously burst: this event is automatically detected by the camera and the linear motor the lowers the frame back in the bath. After a short delay of 5 s, a new experiment can start.
@@
The solutions used to generate soap films were prepared with deionized water, adding NaCl in amounts ranging from 1 g/L to 147 g/L. A non-ionic surfactant (Triton X-100) was added with a concentration of three times the Critical Micelle Concentration (CMC) to stabilize the film. This surfactant was chosen to minimize the interactions between the dissolved ions and the surfactant. The static surface tension of the various solutions was measured using a Langmuir trough and was found to be essentially independent of the salinity and equal to 32 mN/m (see App. A). At high concentrations, solutions of NaCl have a viscosity up to 2 times that of pure water (Kestin et al.,, 1981), to isolate the effect of viscosity, we therefore also tested solutions of 5 and 10% Glycerol by weight with the same surfactant. A summary of the experimental conditions is provided in Tab 1 and is separated in two parts: in the drainage part, the experiments are repeated five to ten times, and we explore many different conditions. Note that we also performed experiments with a different concentration of Triton X-100 (1 CMC) and with C12E9, another non-ionic surfactant. In the lifetime section, we focus on obtaining robust statistics in the zero salinity case, but changing the relative humidity.
@@
In order to explore the effect of mixing of the air in the chamber, we use Rayleigh-Bénard convection. We use a heating mat at the bottom of the chamber and Peltier plates at the top to achieve the desired temperature difference. The temperatures at the top and bottom surface are measured and controlled using analog sensors (Texas Instruments LM35DZ/LFT4), and the temperature at the level of the soap film is also measured to ensure that it remains between 21 and 22C. The heating mat and Peltier plates are duty-cycle controlled such that the wall temperature is constant. The Rayleigh number controlling the instability is $Ra=\rho_{a}\beta_{a}\Delta T\ell^{3}g/\mu_{a}\alpha_{a}=\mathcal{O}(10^{7})$, with $\ell=60$ cm the height of the chamber, $\beta$ the thermal expansion coefficient, $\rho_{a}$ the density of air, $\mu_{a}$ the viscosity of air, $\alpha_{a}$ the thermal diffusivity of air, and a temperature difference $\Delta T$. At this Rayleigh number, the flow in the chamber is turbulent (Drazin,, 2002) and the air surrounding the film is therefore homogeneous in space as any slow or local variations of the experimental conditions are suppressed. The instability sets in air in the chamber into motion, and the typical root-mean-square velocity is 5 mm/s (measured using Particle Image Velocimetry). In the case where mixing is not activated, the air in the chamber can still have some motion, but it is much slower: the air inlets are placed at the bottom of the chamber and a diffuser restricts the air inlet flow velocity. A small ($<5$%) but stable vertical humidity gradient forms in the chamber, these effects resulting in a velocity too small to measure with our PIV setup ($<1$ mm/s). We detail the effect of mixing the air in the chamber in Sec. 5; in the drainage section (Sec. 4) there is no mixing, but we show in Fig. 5 that convection does not alter the drainage significantly.
@@ 3 Qualitative observations
From moderate salinities and at any humidity, an instability is occurring near the top of the film. There, thick patches are formed at the interface between the colored film and the black film (brown plumes at the top of Fig. 2 (b) and yellow area in panel (a)). These patches are thicker than the ones surrounding them and therefore fall down in the film because of buoyancy until they reach the level that matches their thickness. This instability is reminiscent of a Rayleigh-Taylor instability, or to the one observed by Shabalina et al., (2019). This instability occurs in many of the cases studied: with other surfactants and in cases where glycerol was used instead of salt. It is therefore not specific to salted films and most likely simply a feature of long-lasting films and was previously studied in different contexts (Seimiya and Seimiya,, 2021).
@@ 4 Drainage
Using the videos of the film draining and the interferometry technique described in Sec. 2 we are able to measure the thickness of any given point, over time. A typical drainage curve for all the cases we explore in shown in Fig. 3 (a): the thickness is a few microns and drains over a timescale ranging from a few tens of seconds to minutes. We can also recover vertical thickness profiles by interpolating between consecutive interference fringes, at a given instant. This technique works from a few seconds after the film has stopped being pulled out ($t\gtrsim 2$ s) and outside the regions where marginal regeneration patches are present ($z\lesssim 15$ mm). This allows us to investigate how the vertical profile of thickness evolves over time (illustrated in Fig. 3 (b)). The thickness increases almost linearly when moving towards the bottom of the film, with a slope $\partial h/\partial z$ that decreases in time.
@@
Within the parameter range tested, film drainage is remarkably similar: without any rescaling the drainage curves almost collapse onto a single curve (Fig. 3 (c)). In particular the fact that the drainage curves for different humidities are the same (different markers) shows that in the regime of interest with films lasting at most a few hundred seconds, evaporation has a negligible role in the thinning. This property (which was previously noted in the case of continuously drawn films Champougny et al., (2018)) holds true regardless of salinity, despite the known effect of salt (or glycerol) to alter evaporation rates (El-Dessouky et al.,, 2002; Roux et al.,, 2022). The effect of salinity (different colors), is to slightly shift the curves without altering the initial thickness of the film. The rate at which the film drains is therefore altered by salinity, but this change remains moderate in the conditions tested. Similarly, increasing the viscosity (orange and pink markers) slightly shifts the curves and slows down drainage. At very long times, the curves corresponding to Glycerol and NaCl solutions start to deviate from each other. We interpret this result as being due to the qualitatively different features visible in the films. The collapse of the drainage is robust to a change of surfactant and / or surfactant concentration in the cases we explored (not shown on Fig. 3).
At long times (after about 10 s), the drainage curve exhibits a power law behavior in time with an exponent close to -1 (i.e. $h\propto 1/t$). Monier et al., (2024) recently proposed a general form for the evolution of thickness at a given height $z^{*}$ in the film: $h(z^{*},t)\propto(\kappa/(t-t_{0}))^{m}$ with $\kappa$, $t_{0}$ and $m$ constants that can vary slightly depending on the experiment. Using data from the literature as well as a series of experiment with various viscosities and frame geometries, they found an exponent $m$ between 1 and 2, with an average value of 1.41. By fitting the same law onto our present experiment, we find an exponent $m=1.3\pm 0.2$. As discussed in their work, we also find a small negative value for $t_{0}$ corresponding to the initial finite thickness of the films. The exponent found here is coherent with previous experiments on soap films where evaporation does not play a major role and where the main draining mechanism is the exchange of patches of fluid with the Plateau borders at the sides of the film (Berg et al.,, 2005). These films are said to be mobile (Mysels,, 1964) in opposition to rigid films where the interface is contaminated to the point where it applies a no-slip boundary condition to the interstitial fluid. In that case, the equation describing the thinning of the film is (Mysels,, 1959): $\partial h/\partial t=-\rho g/4\mu\times h^{2}\partial h/\partial z$, with $\mu$ and $\rho$ the viscosity and density of the liquid, respectively. A downwards Poiseuille flow develops inside of the film, leading to a thickness varying like $h\sim\sqrt{z/t}$, clearly not compatible with the present experiments. The surfactant gradient stabilizing the film can instead be taken into account by introducing a finite stress at the interface. This stress a priori depends on the surfactant used and on the surfactant concentration and should be computed by introducing the equation of state for the surface excess of surfactant (Champougny et al.,, 2015). Under the simplifying assumption of a constant slip length (or stress) at the interface, previous authors have numerically recovered a thinning law close to $t^{-1}$ (Berg et al.,, 2005) with an exponent that depends on the choice of boundary condition. The variations of the drainage exponent $m$ visible in the experiment can therefore be interpreted as a change of boundary condition for different cases.
@@
Figure 4: (a) Dimensionless thickness as a function of $z/H$, color indicates time.  (b) Thickness as a function of the rescaled parameter following Eq. 1 The data shown in (a) and (b) corresponds to a film with a salinity of 72 g/L and a relative humidity of 70 %. The black line is a linear fit with slope 2.0.  (c) Thickness as a function of the rescaled parameter, for all salinities and relative humidities tested. Colors and markers are identical to Fig. 3: markers indicate relative humidity and colors indicate salinity or Glycerol concentration. Each marker corresponds to an average between 5 and 10 repetitions. (Inset) Measured proportionality constant as a function of the salinity (Eq. (1)). Errorbars represent the dispersion between repetitions.
@@
with $h_{\rm ref}=7$ µm, the thickness at generation and $t_{\nu}$ the viscous drainage time. This typical timescale can be obtained by considering the thin film equation (Mysels,, 1959), giving:
$t_{\nu}=\frac{\mu H}{\rho gh_{\rm ref}^{2}}.$ (2)
@@
For the solutions considered here this typical viscous time is 40 to 50 s depending on the salinity. Using the drainage scaling in combination with the viscous timescale, we can see that the thickness data for all solutions reasonably collapses onto a straight line (Fig. 4 (c)). The slope $\mathcal{A}$ of this linear relationship is an order $\mathcal{O}(1)$ non-dimensional constant, which can be evaluated and is independent on the salinity or on the viscosity ($\mathcal{A}\approx 0.3$, inset). Using glycerol or NaCl, the effect of the viscosity is therefore fully accounted for through the viscous timescale $t_{\nu}$.
There is a small visible offset between different solutions that may be attributed to the drainage at a fixed height following a power law close to but not exactly minus one or the shape vertical profile not being always perfectly linear. Using the fitted exponent $m$ and the more general expression proposed by Monier et al., (2024) to rescale the data yields a very similar result to the one plotted on Fig 4 (c), but our approach presents the advantage of allowing us to measure the slope $\mathcal{A}$ without any fitting parameters, and this single parameter is an order 1 constant.
@@
The drainage scaling presented here is robust for all conditions tested of salinity and relative humidity but also to a change of surfactant concentration or surfactant (Triton X-100 at 1 CMC or C12E9 at 3 CMC). This law is only valid from a few seconds after the start of the drainage (Fig. 3 (c)) and applies only in the bulk of the film, away from edge effects: marginal regeneration at the bottom or sides of the film and the boundary between the colored and black film at the top. As noted above, if the film was rigid instead of mobile (following the distinction from Mysels, (1964)), the thickness should follow $h\propto\sqrt{z/t}$ and this drainage scaling would break down. For solutions with moderate salinities (up to seawater concentration), this study of film drainage therefore allows us to conclude that (i) the only effect of NaCl on thin film drainage is through viscosity and (ii) relative humidity (and therefore evaporation) does not play a significant role in the drainage, regardless of salinity.
@@
This last remark should be contrasted with the very strong effect that relative humidity has on film lifetime. For a given solution, the lifetime can range from only a few seconds at 30% humidity to tens of minutes in an almost saturated atmosphere at 90% humidity. In the following we therefore focus on the effect of relative humidity on the lifetime of the films and obtain a scaling law explaining this dependence.